OpenStudy (anonymous):

A street light is mounted at the top of a 15-ft-tall pole. A man 6 ft tall walks away from the pole with a speed of 7 ft/s along a straight path. How fast is the tip of his shadow moving when he is 30 ft from the pole?

OpenStudy (amistre64):

looks like you want to set up a relationship associated with right tirangles, or similar triangles

OpenStudy (amistre64):



OpenStudy (amistre64):

l/m = s/(d+s) l(d+s) = ms l'(d+s) + l(d+s)' = m's + ms' since l and m are not changeing, l' and m'=0 l(d+s)' = ms' solve for s' to determine d'+s'

OpenStudy (anonymous):

thank you i had already figured it out. the answer was 35/3 ft/s